### v1

# secReferences

Abstract
The kidney paired donation (KPD) program provides an innovative solution to overcome incompatibility challenges in kidney transplants by matching incompatible donor-patient pairs and facilitating kidney exchanges. To address unequal access to transplant opportunities, there are two widely used fairness criteria: group fairness and individual fairness. However, these criteria do not consider protected patient features, which refer to characteristics legally or ethically recognized as needing protection from discrimination, such as race and gender. Motivated by the calibration principle in machine learning, we introduce a new fairness criterion: the matching outcome should be conditionally independent of the protected feature, given the sensitization level. We integrate this fairness criterion as a constraint within the KPD optimization framework and propose a computationally efficient solution. Theoretically, we analyze the associated price of fairness using random graph models. Empirically, we compare our fairness criterion with group fairness and individual fairness through both simulations and a real-data example.
Keywords: calibration, integer programming, kidney paired donation, price of fairness, random graph.
Introduction
1.1 Kidney paired donation programs
Kidney transplantation is the preferred treatment for end-stage renal disease (ESRD), offering significant improvements in both quality of life and survival compared to dialysis. However, a major obstacle is the incompatibility between donors and patients, often due to mismatches in blood type or human leukocyte antigens (HLA). According to the United Network for Organ Sharing (UNOS) and the Organ Procurement and Transplantation Network (OPTN), over 90,000 patients were on the kidney transplant waiting list at the end of 2023.
To overcome these incompatibility challenges, kidney paired donation (KPD) programs have been developed as an innovative solution. These programs match incompatible donor-patient pairs using a virtual crossmatch, a preliminary compatibility test, and facilitate kidney exchanges, allowing patients to receive kidneys from compatible donors through mutual exchange. These exchanges are called exchange cycles or simply cycles, with formal definitions provided in Section 2. The primary goal of KPD programs is to maximize the number of successful transplants or optimize generalized utilities based on predicted transplantation or survival outcomes.
Despite their promise, planned cycles may fail for various reasons, such as illness, pregnancy, or the death of a patient or donor, scheduling conflicts, or discrepancies between virtual and laboratory crossmatch results. These uncertainties necessitate the consideration of recourse strategies, which identify alternative transplant opportunities within the original cycle. Even when an entire cycle cannot proceed, smaller unaffected sub-cycles may still be viable. By accounting for these uncertainties, KPD programs can further maximize the expected number of successful transplants or optimize the expectation of some general utilities.
Following the framework of Klimentova et al. (2016), KPD programs adopt three main recourse strategies. The first is no recourse, where a cycle either proceeds fully or fails entirely (e.g. Li et al. 2014). The second, internal recourse, identifies the sub-cycle with the highest utility among those unaffected by the failure (e.g. Pedroso 2014). The third, subset recourse, considers broader subsets that may include multiple cycles, enabling alternative arrangements when uncertainties arise (e.g. Bray et al. 2018). Each strategy computes expected utilities accordingly, with KPD programs aiming to maximize these expected utilities.
In addition to recourse strategies, we can integrate uncertainties into the optimization framework through other approaches, such as look-ahead strategy (Wang et al. 2017) and robust optimization (McElfresh et al. 2019).
1.2 Fairness concerns in KPD programs
Despite the success of KPD programs, important fairness concerns arise, particularly regarding unequal access to transplant opportunities. These disparities stem from two main factors. The first factor is differences in patients’ HLA sensitization levels. A patient’s HLA sensitization level is measured by their panel-reactive antibody (PRA) score, which reflects the likelihood of HLA incompatibility with a random donor. Patients with high PRA scores, referred to as highly sensitized patients, are more difficult to match and, therefore, have fewer transplant opportunities. In contrast, patients with low PRA scores, known as lowly-sensitized patients, are easier to match and benefit from more transplant opportunities. The second factor is asymmetric blood type compatibility. According to standard ABO blood type compatibility rules, patients with blood type O are harder to match because they can only receive kidneys from O-type donors. Conversely, patients with blood type AB are the easiest to match, as they can receive kidneys from donors of any blood type.
To address these disparities, KPD programs can incorporate fairness constraints to reduce unfairness caused by differences in HLA sensitization and blood type. Two widely used fairness criteria in KPD programs are group fairness and individual fairness. Group fairness focuses on ensuring that highly-sensitized patients receive equitable consideration relative to lowly-sensitized patients (Dickerson et al. 2014; McElfresh et al. 2019; Freedman et al. 2020). In contrast, individual fairness aims to provide balanced selection chances for each patient, ensuring that no one is unfairly disadvantaged (Farnadi et al. 2021). While these two fairness criteria are central to addressing disparities in kidney exchange, other approaches have also been explored. For example, St-Arnaud et al. (2022) incorporate the Nash standard of comparison (or proportional fairness) and Rawlsian justice principles. Ashlagi and Roth (2014), Klimentova et al. (2021), and Carvalho and Lodi (2023) draw on game theory to address fairness within utility-maximization frameworks, considering the interests of stakeholders such as hospitals and regions.
In the context of fair machine learning, a protected feature (or sensitive attribute) refers to a characteristic legally or ethically recognized as needing protection from discrimination. Our key question is how to establish a fairness criterion that ensures equal access to transplant opportunities across patient groups defined by protected characteristics—and how to achieve this in practice. This specific focus has not yet been explored in the KPD literature, as existing fairness criteria do not account for protected patient features.
Some protected features, such as race and gender, are associated with differences in sensitization levels and blood types. For instance, studies have shown that parous women are more likely to develop high sensitization to HLA antigens (Bromberger et al. 2017), making them less compatible with most donors in a KPD program and harder to match. This leads to unequal access to transplant opportunities between females and males. A simplistic approach might aim to balance overall selection rates between genders, but this could inadvertently disadvantage highly-sensitized male patients, as a higher number of highly-sensitized female patients would need to be matched to achieve gender balance. A more equitable approach would balance selection rates within subgroups, such as highly-sensitized females versus males and lowly-sensitized females versus males.
Motivated by this example, we propose a new fairness criterion: the matching outcome should be conditionally independent of the protected feature, given the sensitization level. The randomness associated with this fairness criterion is determined by a randomization policy, as proposed for individual fairness in kidney exchange (Farnadi et al. 2021; St-Arnaud et al. 2022) and general matching problems (García-Soriano and Bonchi 2020; Karni et al. 2022). This approach provides guarantees for average selection rates within protected groups across each sensitization level.
1.3 Fairness in general decision-making problems
Our fairness criterion in KPD programs is defined based on the conditional outcome given protected features, drawing on similar concepts from the literature on general decision-making problems.
Demographic parity ensures fairness by requiring that the rate of positive decisions is consistent across groups defined by protected features, promoting equality in outcomes regardless of group membership. Equalized odds, introduced by Hardt et al. (2016), aligns predictive performance such that the false positive and false negative rates are similar across groups, leading to a fair distribution of errors. Predictive parity, discussed by Chouldechova (2017), ensures fairness by equalizing the positive predictive value (PPV) across groups, thereby making positive predictions equally reliable and trustworthy for all groups. Calibration within groups, explored by Kleinberg et al. (2017), requires that individuals with the same predicted probability have consistent actual outcome rates across groups, ensuring well-calibrated predictions.
These fairness concepts emphasize different priorities: overall outcome equality, error distribution, prediction reliability, or probability calibration. They are frequently incorporated as fairness constraints in statistical optimization problems (e.g., Liebl and Reimherr 2023).
Our fairness criterion in KPD programs is closely aligned with calibration within groups; see a more detailed discussion in Section 3.1.
1.4 Our contributions
This paper makes several contributions to the field of kidney exchange and fairness in allocation. First, we propose a new fairness criterion based on a protected feature, which has not been explored in the kidney exchange literature. We integrate this fairness criterion as a constraint within the optimization framework commonly used in kidney paired donation (KPD) programs. This flexible structure can accommodate other fairness criteria and potential recourse strategies. Furthermore, we propose a computationally efficient solution to the resulting optimization problem.
Second, we investigate the price of fairness associated with our proposed criterion, defined as the relative loss in system efficiency when a fair allocation is prioritized over an optimal (unconstrained) allocation (Bertsimas et al. 2011). Theoretically, we derive an upper bound on the asymptotic price of fairness using random graph models that incorporate ABO blood type distributions. Empirically, through simulation studies, we show that the efficiency loss from implementing our fairness criterion is relatively low.
Our findings align with prior studies on the tradeoff between efficiency and fairness in resource allocation. For example, Dickerson et al. (2014) examine the efficiency loss associated with group fairness in kidney exchange, while Ashlagi and Roth (2014) analyze the price of ensuring individual rationality in multi-hospital kidney exchanges, both employing random graph models with ABO blood types. St-Arnaud et al. (2022) utilize the Nash Social Welfare Program to address the tradeoff between fairness and efficiency. Similarly, Viviano and Bradic (2024) propose a framework for fair policy targeting that balances fairness and efficiency using Pareto optimal treatment allocation rules, offering theoretical guarantees and practical solutions applicable to social welfare contexts.
Review of KPD program in an optimization framework
In this section, we review the notations, terminologies, and optimization framework used in KPD programs. Section 2.1 focuses on the classical optimization problem without incorporating fairness constraints, while Section 2.2 reviews group fairness and individual fairness, introducing an additional constraint to integrate fairness into the framework.
2.1 KPD program without fairness
2.1.1 Classical formulation
We represent a KPD program as a directed graph $G=(V,E)$, where $V=\{v_{1},...,v_{N}\}$ denotes the vertex set and $E$ denotes the edge set. The vertex set $V$ is the set of $N$ incompatible donor-patient pairs. The edge set $E$ consists of all ordered pairs $(v_{i},v_{j})$ that the donor in vertex $v_{i}\in V$ is compatible with the patient in vertex $v_{j}\in V_{p}$. An exchange cycle, or simply cycle, is defined as a sequence of distinct incompatible donor-patient pairs. We denote a cycle $c$ as an ordered sequence of vertices $[c_{1},...,c_{|c|}]$ in $V$, where $|c|$ is the cycle length of $c$, satisfying that $(c_{i},c_{i+1})\in E$ for $11\leq i\leq|c|-$ 1, and $(c_{|c|},c_{1})\in E$. To execute the cycle, the patient in $c_{i+1}$ will receive the donor kidney of $c_{i}$ for $11\leq i\leq k-$ 1, and the patient $c_{1}$ will receive the donor kidney of $c_{k}$. Figure 1 illustrates how a two-way cycle and a three-way cycle work in KPD programs. An exchange plan is a collection of vertex-disjoint cycles in the graph.
donor Adonor Bpatient Apatient B
(a) A two-way cycle [A,B]
donor Adonor Bdonor Cpatient Apatient Bpatient C
(b) A three-way cycle [A,B,C]
Figure 1: Illustration of exchange cycles shown with solid arrows. Transplantations along the dashed arrows cannot proceed due to incompatibility.
In fielded KPD programs, we usually restrict cycles to be no longer than 3. We use $\mathcal{C}$ to denote the set of all cycles under the restrictions on the length of the cycles. We assign a utility $u_{v_{i},v_{j}}$ to each edge $(v_{i},v_{j})\in E$. Based on the edge utility, we define the cycle utility $u(c)=\sum_{i=1}^{|c|-1}u_{c_{i},c_{i+1}}+u_{c_{|c|},c_{1}}$. In a classical KPD program, we aim to select disjoint packing of cycles in $\mathcal{C}$ with the maximum sum of utilities, which can be formulated as the following integer programming:
$\max_{x_{c}\in\{0,1\}}\quad\sum_{c\in\mathcal{C}}x_{c}u(c)\quad\text{subject to}\quad\sum_{c\in\mathcal{C}}x_{c}1_{v\in c}\leq 1,\ \forall\ v\in V.$ (1)
The binary variable $x_{c}$ indicates whether $c$ is selected or not. The objective function in (1) is the sum of utilities of the selected cycles. The constraint in (1) requires that the selected cycles are vertex-disjoint.
2.1.2 Extension: incorporating recourse strategies
To account for uncertainty, we consider the expected utilities in (1), depending on the chosen recourse strategy. Different strategies lead to different calculations of expected utilities.
No-recourse strategy.
Under the no-recourse strategy, the expected utility is determined without adapting to failures. Given the failure probabilities of each vertex in $V$ and each edge in $E$, we can compute the probability of each cycle in $\mathcal{C}$ being executable, where all vertices and all edges in the cycle do not fail. An explicit formula is available to compute the expected utility for this strategy (Li et al. 2014; Klimentova et al. 2016).
Internal-recourse strategy.
The internal-recourse strategy considers adapting to failures within a given cycle. For a cycle $c$, let $\mathcal{M}(c)$ represent the set of all exchange plans involving vertices in $c$. Without loss of generality, assume the elements of $\mathcal{M}(c)$ are ordered by descending utility. Specifically, the utility of $M_{i}(c)$, the $i$-th element in $\mathcal{M}(c)$, is no less than that of $M_{j}(c)$, the $j$-th element, for $i<j$. If all vertices and edges in $M_{1}(c)$ do not fail, $M_{1}(c)$ will be executed. If any vertices or edges in $M_{k}(c)$ fail but those in $M_{k+1}(c)$ do not fail, $M_{k+1}(c)$ will be executed ($k\geq 1$). While there is no explicit formula for this strategy, efficient algorithms are available for computing the expected utility (Li et al. 2014; Pedroso 2014; Klimentova et al. 2016).
Subset-recourse strategy.
The subset-recourse strategy expands the search space to disjoint relevant subsets rather than disjoint cycles, providing more flexibility in adapting to uncertainty. Formally, a relevant subset of size $(k,q)$ is the set of at most $(k+q)$ vertices in graph $G$ inducing a strongly connected subgraph, such that any edge of the paths that provide the strong connectivity belongs to some cycle of size at most $k$, whose vertices are in the subset. We follow the same approach as the internal-recourse strategy to compute the expected utility of each relevant subset. To solve the optimization problem in (1), we must enumerate the new set $\mathcal{C}$, with algorithms given in Klimentova et al. (2016) and Wang et al. (2019).
2.2 KPD program with fairness constraint
2.2.1 Group fairness
The group fairness aims to ensure fairness towards highly-sensitized patients. We partition the vertex set $V$ into $V_{h}\cup V_{e}$, where $V_{h}$ denotes the set of incompatible donor-patient pairs with highly-sensitized patients, and $V_{e}$ denotes the set of incompatible donor-patient pairs with lowly-sensitized patients. Following Dickerson et al. (2014), we formulate the group fairness as a constraint in the optimization problem below
$\max_{x_{c}\in\{0,1\}}\ \sum_{c\in\mathcal{C}}x_{c}u(c)\quad\text{subject to}\quad\sum_{v\in V_{h}}\sum_{c\in\mathcal{C}}x_{c}1_{v\in c}\geq\alpha,\ \sum_{c\in\mathcal{C}}x_{c}1_{v\in c}\leq 1,\ \forall\ v\in V.$ (2)
Specifically, we consider a fairness constraint that the number of highly-sensitized patients involved in the matching is no less than some threshold $\alpha$.
2.2.2 Individual fairness
The individual fairness utilizes a randomization policy to ensure that every patient has a similar chance to be matched. Let $\mathcal{F}\subseteq 2^{\mathcal{C}}$ be the set of possible exchange plans that consist of disjoint cycles in $\mathcal{C}$. We assign a probability distribution $\delta$ on $\mathcal{F}$, which indicates the probability of selecting each exchange plan in $\mathcal{F}$. Let $x_{c}(\delta)$ be the probability that the cycle $c\in\mathcal{C}$ is selected as one of the disjoints subsets. Based on $x_{c}(\delta)$, we can compute the probability of each vertex $v$ to be selected $\delta_{v}=\sum_{c\in\mathcal{C}}x_{c}(\delta)1_{v\in c}$, which is a central quantity in defining individual fairness and our new fairness in Section 3. Recall that $x_{c}$ in (1) and (2) is a binary variable indicating whether $c$ is selected or not, but here $x_{c}(\delta)$ is a continuous variable bounded between 0 and 1. Following Farnadi et al. (2021), We formulate the individual fairness as a constraint in the optimization problem below
$\max_{\delta}\quad\sum_{c\in\mathcal{C}}x_{c}(\delta)u(c)\quad\text{subject to}\quad\sum_{v\in V}|\delta_{v}-\overline{\delta}|^{p}\geq\beta^{p}.$ (3)
Specifically, the individual fairness promotes a similar chance of being selected for each patient. We choose the $L_{p}$ norm of the vector $(\delta_{v}-\overline{\delta})_{v\in V}$ to measure the variation of the selection probability among all the patients, where $\overline{\delta}$ is the average selection probability of all patients. We consider a fairness constraint that the $L_{p}$ norm of the vector $(\delta_{v}-\overline{\delta})_{v\in V}$ is no greater than some threshold $\beta$.
A new fairness criterion based on a protected feature
3.1 Fairness formulation and algorithm
In this section, we introduce a new fairness criterion based on a protected feature $A$ of patients in $V$, motivated by the concept of calibration within group in the context of machine learning. In binary classification, a score function $R(X)$ satisfies calibration within groups if $\mathrm{Pr}(Y=1\mid R(X)=r,A=a)=r$ for all score values $r$ and group level $a$, where $Y$ is the outcome variable and $X$ is the feature variables. A slightly weaker definition (Corbett-Davies et al. 2023) only requires that $Y$ is conditionally independent of $A$ given the score $R(X)$. Our fairness definition is based on this weaker definition. In KPD programs, we view $Y$ as the selection indicator and $R$ as the sensitization level of a patient. Therefore, we define our new fairness as satisfying that the selection indicator is conditionally independent of the protected feature given the sensitization level. In other words, at each sensitization level, the matching outcome is independent of the protected feature. We can view this fairness criterion in KPD programs as a reverse problem of that in machine learning. In machine learning, the randomness is due to the underlying population model, and our goal is to construct a score $R(X)$ satisfying the fairness condition. In KPD programs, we need to determine the randomness such that the observed sensitization level satisfies the fairness condition.
For simplicity of presentation, we assume the protected feature $A$ is binary with two levels $\{0,1\}$ and the sensitization $R$ has $M$ levels $\{r_{1},r_{2},...,r_{M}\}$. In practice, it is common that $M=2$ where we partition all patients into highly-sensitized and lowly-sensitized patients, or $M=3$ where we partition all patients into highly-sensitized, moderately-sensitized and lowly-sensitized patients. The joint of variables $A$ and $R$ partitions all patients into $2M$ subgroups, denoted as $\{V_{ij}\}_{0\leq i\leq 1,1\leq j\leq M}$, where $V_{ij}=\{v\in V:A(v)=i,R(v)=r_{j}\}$. As in Section 2.2.2, we assign a probability distribution $\delta$ on the set of exchange plans $\mathcal{F}$, which indicates the probability of selecting each exchange plan in $\mathcal{F}$. Moreover, $\delta_{v}$ is the probability of each vertex $v$ to be selected in the end, and thus $V_{ij}^{-1}\sum_{v\in V_{ij}}$ is average selection rate in $V_{ij}$. Given the sensitization level $R=r_{j}$, our fairness criterion restricts the average selection rates in $V_{0j}$ and $V_{1j}$ to be close. It is natural to impose constraints that the absolute difference in these average selection rates is bound by a constant $l_{j}$. Therefore, we formulate the problem under our fairness criterion as
$\max_{\delta}\ \sum_{c\in\mathcal{C}}x_{c}(\delta)u(c)\quad\text{subject to}\ \left|V_{0j}^{-1}\sum_{v\in V_{0j}}\delta_{v}-V_{1j}^{-1}\sum_{v\in V_{1j}}\delta_{v}\right|\leq l_{j},\ \forall 1\leq j\leq M.$ (4)
However, the optimization problem (4) can be difficult to solve because $\mathcal{F}$ is too large to enumerate, let alone the choice of $\delta$. We provide an approach to solve (4) based on Proposition 1 below.
Proposition 1.
There exists a solution to (4) satisfying that at most $M+1$ exchange plans in $\mathcal{F}$ have nonzero selection probability.
Instead of solving the harder optimization problem (4), by Proposition 1, we can solve an equivalent but easier mixed-integer programming:
$\max_{x_{t,c}\in\{0,1\},p_{t}\geq 0}\quad\sum_{t=1}^{M+1}\sum_{c\in\mathcal{C}}p_{t}x_{t,c}u(c)$ (5)
subject $\text{subject to}\quad&\sum_{t=1}^{M+1}p_{t}=1,&&\\ &\sum_{c\in\mathcal{C}}x_{t,c}1_{v\in c}\leq 1,\quad\forall\ v\in V,\ 1\leq t\leq M+1,\\ &|\overline{q}_{j1}-\overline{q}_{j1}|\leq l_{j},\quad\forall 1\leq j\leq M,&&\\ \end{aligned}$ (6)
where we define $\overline{q}_{ij}=|V_{ij}|^{-1}\sum_{v\in V_{ij}}\sum_{t=1}^{M+1}\sum_{c\in\mathcal{C}}p_{t}x_{t,c}1_{v\in c}$ as the average selection rate in $V_{ij}$. The optimization problem (5)–(6) involves $(M+1)|\mathcal{C}|$ binary variables and $M+1$ continuous variables. It can be efficiently solved by various optimizers, e.g., Gurobi Optimizer.
For whichever choice of parameters $l_{1},...,l_{M}\geq 0$, there always exists a solution to (5)–(6). The strength of the fairness constraint depends on the parameters $l_{1},...l_{M}$. In practice, we suggest two candidate values for $l$’s: $l_{j}=1/\min\{|V_{1j}|,|V_{2j}|\}$ and $l_{j}=1/\max\{|V_{1j}|,|V_{2j}|\}$. We can interpret these two candidate values as the desired precision based on the larger or smaller subgroup. The former represents a weaker fairness constraint, and the latter represents a stronger one.
3.2 Prediction of individual selection probability
Selection probability $\delta_{v}$ is a central quantity in defining both individual fairness and our new fairness. Since a KPD program is not static but dynamic, a natural statistical question to ask is how to predict individual selection probabilities before more incompatible donor-patient pairs enter the pool for the next round of exchange allocation. We provide a solution based on sample splitting. The following discussion can accommodate any fairness criteria within our general framework.
Our method utilizes historical data of incompatible donor-patient pairs independent of the current and future pairs. Suppose the historical pool consists of incompatible donor-patient pairs denoted as $\{\tilde{v}_{1},...,\tilde{v}_{N_{0}}\}$, and the current pool consists of pairs denoted as $\{v_{1},...,v_{N_{1}}\}$. Assume that the exchange allocation occurs when the size of the incompatible pairs pool reaches $N$, where $N_{1}<N<N_{0}+N_{1}$. The prediction procedure can be described in Algorithm 1 below.
Algorithm 1 Prediction procedure with selection probabilities
1:  Input: Historical pool of vertex set $\{\tilde{v}_{1},\dots,\tilde{v}_{N_{0}}\}$, current pool of vertex set $\{v_{1},\dots,v_{N_{1}}\}$, and number of repetitions $B$.
2:  Output: Prediction of $\delta_{v}$ for each $v\in\{v_{1},\dots,v_{N_{1}}\}$.
3:  for $b=1$ to $B$ do
4:     Sample $\{\tilde{v}_{1}^{(b)},\dots,\tilde{v}_{N-N_{1}}^{(b)}\}$ from $\{\tilde{v}_{1},\dots,\tilde{v}_{N_{0}}\}$ without replacement.
5:     Determine the edge set $E^{(b)}$ in $V^{(b)}=\{v_{1},\dots,v_{N_{1}},\tilde{v}_{1}^{(b)},\dots,\tilde{v}_{N-N_{1}}^{(b)}\}$.
6:     Based on the graph $(V^{(b)},E^{(b)})$, solve (4) to obtain the selection probability of $\delta_{v}^{(b)}$ for each $v\in\{v_{1},\dots,v_{N_{1}}\}$.
7:  end for
8:  Return: mean and quantiles of $\{\delta_{v}^{(1)},...,\delta_{v}^{(B)}\}$ as the mean prediction and interval prediction of $\delta_{v}$, for each $v\in\{v_{1},\dots,v_{N_{1}}\}$.
In practice, the computational complexity could be very high to enumerate the cycles or relevant subsets of $\{v_{1},...,v_{N_{1}},\tilde{v}_{1}^{*},...,\tilde{v}_{N-N_{1}}^{*}\}$ for $B$, say 1000, times. When $M$ is small, it could be more computationally efficient to first enumerate the cycles or relevant subsets $\{v_{1},...,v_{N_{1}},\tilde{v}_{1}^{*},...,\tilde{v}_{N_{0}}^{*}\}$, and then filter the cycles or relevant subsets with vertices in $\{v_{1},...,v_{N_{1}},\tilde{v}_{1}^{*},...,\tilde{v}_{N-N_{1}}^{*}\}$ for each replication. However, when $N_{0}$ is extremely large, it could be impossible to enumerate the cycles or relevant subsets $\{v_{1},...,v_{N_{1}},\tilde{v}_{1}^{*},...,\tilde{v}_{N_{0}}^{*}\}$. One solution is to split $\{\tilde{v}_{1},...,\tilde{v}_{N_{0}}\}$ into disjoint subsets of appropriate size and to enumerate the cycles or relevant subsets within each subset accordingly.
Price of the new fairness criterion under random graph models
In this section, we establish theoretical guarantees for the price of fairness associated with our new fairness criterion using a random graph model that incorporates ABO blood types. We describe the model assumptions below.
Recall that a donor and a patient are compatible if they match in both blood type and HLA. We assume blood type compatibility follows standard medical guidelines: AB patients can receive kidneys from donors of any blood type, A and B patients can receive from donors of their own type or type O, while O patients can only receive from type O donors. We assume HLA compatibility follows binomial distributions. Specifically, we randomly assign each patient a PRA score, representing the probability of being HLA incompatible with any donor. We assume the PRA scores can take discrete values $\{r_{1},\dots,r_{M}\}$, which also determine sensitization levels $\{r_{1},\dots,r_{M}\}$; and we assume HLA compatibility between different donor-patient pairs is independent. The vertex set $V$ is formed by independently drawing donor-patient pairs from an underlying population, keeping only incompatible pairs until a total of $N$ incompatible pairs is reached. The edge set $E$ is determined by the compatibility between donor-patient pairs in $V$. Regarding the utility assignment, we assume $u(c)$ only depends on the induced subgraph of vertices in $c$, independent of the protected feature of vertices in $c$.
Similar to group fairness that prioritizes highly-sensitized patients, our new fairness criterion needs to prioritize some subgroups of patients based on the protected feature, and further balance their average selection probabilities. In Section 4.1, we derive a general result quantifying the efficiency loss due to subgroup prioritization in KPD programs. Specifically, we focus on prioritizing patients with either $A=1$ or $A=0$, given their blood type and sensitization level. In Section 4.2, we apply the result to establish theoretical guarantees for the price of fairness.
4.1 Efficiency loss of subgroup prioritization
Recall the definition of $V_{ij}$ in Section 3.1. We further write $V_{ij}$ as the union
$V_{ij}=\cup_{b_{1},b_{2}\in\{O,A,B,AB\}}\{V_{b_{1},b_{2},i,j}\},$
where $V_{b_{1},b_{2},i,j}$ is the subset of vertices in $V_{ij}$ with donor blood type $b_{1}$ and patient blood type $b_{2}$. For random graph $G$, consider again the optimization problem in 1:
$\max_{x_{c}\in\{0,1\}}\quad\sum_{c\in\mathcal{C}}x_{c}u(c)\quad\text{subject to}\quad\sum_{c\in\mathcal{C}}x_{c}1_{v\in c}\leq 1,\ \forall\ v\in V$ (7)
but here, we allow for general utility $u$ and a set of cycles or relevant subsets $\mathcal{C}$ with some length limits. Let $\mathcal{P}$ denote the set of indices $(b_{1},b_{2},i,j)$, where the subgroup $V_{b_{1},b_{2},i,j}$ should be prioritized over $V_{b_{1},b_{2},1-i,j}$. We consider the optimization problem under the constraint of prioritizing these subgroups in $\mathcal{P}$:
$\max_{x_{c}\in\{0,1\}}\quad\sum_{c\in\mathcal{C}}x_{c}u(c)$ (8)
subject $\text{subject to}&\quad\sum_{c\in\mathcal{C}}x_{c}1_{v\in c}\leq 1,\ \forall\ v\in V,\\ &\left\{\sum_{v\in V_{b_{1},b_{2},i,j}}\sum_{c\in\mathcal{C}}x_{c}1_{v\in c}-|V_{b_{1},b_{2},i,j}|\right\}\cdot\left\{\sum_{v\in V_{b_{1},b_{2},1-i,j}}\sum_{c\in\mathcal{C}}x_{c}1_{v\in c}\right\}=0,\ \forall(b_{1},b_{2},i,j)\in\mathcal{P}.\end{aligned}$ (9)
The constraint (9) implies that no patients in $V_{b_{1},b_{2},1-i,j}$ can be matched or all patients in $V_{b_{1},b_{2},i,j}$ must be matched, if the subgroup $V_{b_{1},b_{2},i,j}$ is prioritized over $V_{b_{1},b_{2},1-i,j}$. The following Proposition 2 is the main result of this subsection.
Proposition 2.
For a random graph with size $N$, the difference between the maximums achieved in the optimization problem (7) and the optimization problem (8)–(9) is $o(N)$, almost surely as $N\to\infty$.
Proposition 2 shows that we can prioritize patients with either $A=1$ or $A=0$, given their blood type and sensitization level, with ignorable relative efficiency loss when the random graph is large. The result is useful to derive upper bounds for the price of our new fairness, defined as the relative overall utility loss due to the fairness constraint in (4). We present these results in the next subsection.
4.2 Upper bounds for price of the new fairness
4.2.1 Optimizing some general utilities
First, we apply Proposition 2 to the scenario with general utilities, allowing for potential recourse strategies. Let $\mu_{b_{1},b_{2},r,a}$ denote the probability of sampling an incompatible donor-patient pair with donor blood type $b_{1}$, patient blood type $b_{2}$, patient sensitization level $r$, and patient sensitive group level $a$. We can obtain a crude upper bound for the price of fairness in Proposition 3 below.
Proposition 3.
The price of fairness due to the fairness constraint in (4) is no greater than
$\max_{b_{1},b_{2},r}\max\left\{\frac{\mu_{b_{1},b_{2},r,1}\overline{\mu}_{r,0}-\mu_{b_{1},b_{2},r,0}\overline{\mu}_{r,1}}{(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})\overline{\mu}_{r,0}},\frac{\mu_{b_{1},b_{2},r,0}\overline{\mu}_{r,1}-\mu_{b_{1},b_{2},r,1}\overline{\mu}_{r,0}}{(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})\overline{\mu}_{r,1}}\right\}$
almost surely as $N\to\infty$, where $\overline{\mu}_{r,1}=\sum_{b_{1},b_{2}}\mu_{b_{1},b_{2},r,1}$ and $\overline{\mu}_{r,0}=\sum_{b_{1},b_{2}}\mu_{b_{1},b_{2},r,0}$.
If the blood type distributions are balanced across all subgroups defined by different levels of $A$ and $R$, i.e. $\mu_{b_{1},b_{2},r,0}/\overline{\mu}_{r,0}=\mu_{b_{1},b_{2},r,1}/\overline{\mu}_{r,1}$ for all $b_{1},b_{2},r$, then the upper bound in Proposition 3 is 0. If the blood type distributions are not balanced within a specific subgroup level $A=a$ and $R=r$, and the optimal solution only matches patients in this subgroup, then the upper bound in Proposition 3 is attainable.
4.2.2 Maximizing the number of transplants without recourse strategies
Then, we apply Proposition 2 to the scenario that maximizes the number of transplants without any recourse strategies. In this scenario, an explicit optimal allocation is explicitly available in Ashlagi and Roth (2011).
We introduce the following assumptions to simplify the decomposition of the four-way probability $\mu_{b_{1},b_{2},r,a}$. First, we assume that the patient and donor in each incompatible pair share the same protected feature level. Second, we assume that the distributions of sensitization levels are consistent across protected feature levels. While these assumptions are not strictly necessary, they facilitate the decomposition of $\mu_{b_{1},b_{2},r,a}$ into more manageable terms.
Specifically, let $\mu_{a}$ represent the frequency probability of the protected feature level $A=a$. Define $\mu_{O\mid a}$, $\mu_{A\mid a}$, $\mu_{B\mid a}$, and $\mu_{AB\mid a}$ as the frequency probabilities of blood types O, A, B, and AB, respectively, within the protected feature level $A=a$. Under these assumptions, there exists a constant $c$ such that $\mu_{b_{1},b_{2},r,a}=cr\mu_{a}\mu_{b_{1}\mid a}\mu_{b_{2}\mid a}$ for all $b_{1},b_{2}\in{O,A,B,AB}$, $r\in{r_{1},\dots,r_{M}}$, and $a\in{0,1}$.
Moreover, let $\overline{\mu}_{O},\overline{\mu}_{A},\overline{\mu}_{B},\overline{\mu}_{AB}$ denote the frequency probability of blood types O, A, B, AB, respectively, among the whole population. Let $\overline{\gamma}$ be the average PRA score among the whole population. We can obtain a more precise upper bound for the price of fairness in Proposition 4 below.
Proposition 4.
Assume $1.5\overline{\mu}_{A}>\overline{\mu}_{O}>\overline{\mu}_{A}>\overline{\mu}_{B}>\overline{\mu}_{AB}$ and $\overline{\gamma}<0.4$. Let $\phi_{b_{1},b_{2}}=\sum_{k=0}^{1}\mu_{k}\overline{\gamma}\mu_{b_{1}\mid k}\mu_{b_{2}\mid k}$, for $b_{1},b_{2}\in\{O,A,B,AB\}$. Let
$\begin{split}T_{a}(r)=&\mu_{a}r(\mu_{O\mid a}+\mu_{AB\mid a}-\mu_{O\mid a}\mu_{AB\mid a}+\mu_{A\mid a}^{2}+\mu_{B\mid a}^{2})+2\mu_{a}\mu_{A\mid a}\mu_{B\mid a},\\ S_{a}(r)=&T_{a}(r)+\mu_{a}\{\mu_{O\mid a}(1-\mu_{O\mid a})+\mu_{A\mid a}\mu_{AB\mid a}+\mu_{B\mid a}\mu_{AB\mid a}\},\\ Q(r)=&T_{1}(r)+T_{0}(r)+\phi_{B,AB}+\phi_{O,AB}+\phi_{A,AB}+\phi_{A,O}+\phi_{AB,O}+\phi_{O,B},\\ \end{split}$
for $a=0,1$ and $r\in\{r_{1},...,r_{M}\}$, and
$\begin{split}R_{a}=&\min\left\{\phi_{B,AB},2\mu_{a}\mu_{B\mid a}\mu_{AB\mid a}-\phi_{B,AB}\right\}+\min\left\{\phi_{O,AB}+\phi_{A,AB},2\mu_{a}\mu_{A\mid a}\mu_{AB\mid a}-\phi_{O,AB}-\phi_{A,AB}\right\}\\ &+\min\left\{\phi_{A,O}+\phi_{AB,O},2\mu_{a}\mu_{O\mid a}\mu_{A\mid a}-\phi_{A,O}-\phi_{AB,O}\right\}+\min\left\{\phi_{O,B},2\mu_{a}\mu_{O\mid a}\mu_{B\mid a}-\phi_{O,B}\right\},\\ \end{split}$
for $a=0,1$. Then, the price of fairness due to the fairness constraint in (4) is no greater than
$\max_{r\in\{r_{1},...,r_{M}\}}\max\left\{\frac{S_{1}T_{0}-S_{0}T_{1}-S_{0}R_{1}}{S_{1}Q},\frac{S_{0}T_{1}-S_{1}T_{0}-S_{1}R_{0}}{S_{0}Q},0\right\}$
almost surely as $N\to\infty$.
The condition $1.5\overline{\mu}_{A}>\overline{\mu}_{O}>\overline{\mu}_{A}>\overline{\mu}_{B}>\overline{\mu}_{AB}$ gives a mild constraint on the blood types distribution of the whole population. The condition $\overline{\gamma}<0.4$ implies that most patients in the population are not highly-sensitized. Both assumptions are standard and appear in Ashlagi and Roth (2011) and Dickerson et al. (2014). As an application of Proposition 4, we present the following two hypothetical examples to illustrate the asymptotic upper bounds for the price of fairness are very small. The data come from the distribution of blood types in the United States as of 2021, according to the American Red Cross.
As the first example, suppose there are two ethnicity groups, 80% white American and 20% African American, with $\mu_{O\mid 1}=0.45,\mu_{A\mid 1}=0.4,\mu_{B\mid 1}=0.11,\mu_{AB\mid 1}=0.04$ and $\mu_{O\mid 0}=0.51,\mu_{A\mid 0}=0.26,\mu_{B\mid 0}=0.19,\mu_{AB\mid 0}=0.04$. Moreover, there are three sensitization levels $\{0.05,0.45,0.9\}$. Then, with high probability, the price of fairness converges to 0 for any $0.40.05<\overline{\gamma}<$ 0.4. As the second example, suppose there are two ethnicity groups, 90% white American and 10% Asian American, with $\mu_{O\mid 1}=0.45,\mu_{A\mid 1}=0.4,\mu_{B\mid 1}=0.11,\mu_{AB\mid 1}=0.04$ and $\mu_{O\mid 0}=0.4,\mu_{A\mid 0}=0.275,\mu_{B\mid 0}=0.255,\mu_{AB\mid 0}=0.07$. Moreover, there are five sensitization levels $\{0.05,0.25,0.45,0.65,0.9\}$. Then, with high probability, the price of fairness is lower than 0.01 for any $0.090.05<\overline{\gamma}<$ 0.09, and converges to 0 for any $0.400.09\leq\overline{\gamma}<$ 0.40. The two examples show that the price of fairness can be no greater than $1\%$ for large graphs in real practice.
Under a similar random graph model, Dickerson et al. (2014) claim that the price of group fairness is no greater than $2/33$, as $n\to\infty$; Ashlagi and Roth (2011) claim that the relative efficiency loss for individual rationality is only about $1\%$ in multi-hospital kidney exchange. All these results give very low efficiency loss because, in large random graph models, there is a rich set of edges from each vertex such that we can easily adjust for the optimal solution such that the fairness constraint is satisfied. Beyond the kidney exchange setting, Bertsimas et al. (2011) provide an upper bound for the price of proportional fairness and the price of max-min fairness, which are close to 1, in general allocation problems.
Numerical studies
In this section, we present simulation studies based on the random graph models described in Section 4 and real data from the UNOS dataset. All the optimization problems are solved by Gurobi (version 11.0) in R.
5.1 Simulation under the random graph models
We first conduct a simulation study under the random graph models to evaluate the numerical performance of different fairness criteria. We consider a binary protected feature that indicates if one is white or non-white. Specifically, we fix 40 white incompatible pairs (80%) and 10 non-white incompatible pairs (20%). Among the 40 white pairs, 28 patients (70%) are lowly-sensitized, 8 patients (20%) are moderately-sensitized, and 4 patients (10%) are highly-sensitized; and among the 10 non-white pairs, 7 patients (70%) are lowly-sensitized, 2 patients (20%) are moderately-sensitized, and 1 patient (10%) is highly-sensitized. The PRA scores are set to be 0.9, 0.45 and 0.05 for highly-sensitized, moderately-sensitized, and lowly-sensitized patients, respectively. The distribution follows the analysis in Saidman et al. (2006).
For non-white donors and patients, we simulate the blood type from a multinomial distribution (O: 51%, A: 26%, B: 19%, AB: 4%). For white donors and patients, we simulate the blood type from another multinomial distribution (O: 45%, A: 40%, B: 11%, AB: 4%). If a donor and a patient are blood-type compatible, they are incompatible with the probability of the patient’s PRA score; otherwise, they are incompatible with probability 1. Given the patients, we repeat sampling the donors and dropping compatible pairs until the numbers of incompatible pairs are reached. Similarly, based on the blood types and PRA scores, we simulate the edges of the graph. That is, we simulate the compatibility for the donor and patient of every two vertices in the graph. We only allow cycles of length at most 3, and we do not consider any recourse strategies in this section.
We aim to compare the three fairness criteria, group fairness, individual fairness, and our new fairness, with no fairness as a baseline result. We consider the two ways of parameter selection in Section 3.1 for our new fairness. For group fairness, we consider the two ways of parameter selection. The first way is to choose the largest possible $\alpha$ such that (2) is solvable, as discussed in Dickerson et al. (2014). This represents the strongest possible constraint that we must maximize the number of matched highly-sensitized patients. The second one chooses the largest possible $\alpha$ such that the overall utility is equal to that in (1), as discussed in Freedman et al. (2020). In other words, among all the exchange plans that maximize the objective function in (2), we consider the one that maximizes the number of matched highly-sensitized patients. For individual fairness, we choose the $L_{2}$ norm of the vector $(\delta_{v}-\overline{\delta})_{v\in V}$ and two candidate values $\{(0.15/N)^{1/2},(0.25/N)^{1/2}\}$ of $\beta$ in (3). That is, we require the sample variance of $(\delta_{v}-\overline{\delta})_{v\in V}$ to be no greater than $0.15$ and $0.25$, respectively; and the candidate value $(0.15/N)^{1/2}$ represents a stronger constraint and the candidate value $(0.25/N)^{1/2}$ represents a weaker constraint.
Figure 2 presents the average selection rates of different fairness criteria within each subgroup. From Figure 2, group fairness works to increase the selection rates of highly-sensitized patients, and individual fairness works to balance the selection rates of the six subgroups. Differently, our fairness balances the selection rates of the white and non-white patients within each sensitization stratum, instead of the selection rates of all six subgroups. Moreover, our fairness achieves the lowest average absolute difference in selection rates between the two race groups. For the price of fairness, although our fairness (strong) has the lowest average utility, the relative utility loss (28.7-27.91)/28.7=2.8% is relatively low.
Figure 2: Simulation results under random graph models. The average selection rates within each subgroup are calculated over 100 data replications. The error bars represent the mean $\pm$ 1 standard deviation of the absolute differences in selection rates.
5.2 UNOS data analysis
We next conduct a simulation study based on the National UNOS STAR dataset. The National UNOS STAR dataset provides comprehensive transplant records collected by the UNOS, covering donor and recipient characteristics, allocation details, and transplant outcomes. After removing the missing values, the dataset comprises 77,073 records of transplant information for donors and patients, which include details such as blood types, HLA antigens (A1, A2, B1, B2, DR1, DR2), PRA scores, and racial background. The protected feature is set to be a binary variable: white (64.8%) and other racial backgrounds (35.2%). Patients are categorized based on their PRA scores as follows: those with scores above 0.8 are labeled as highly sensitized; those with scores ranging from 0.1 to 0.8 are considered moderately sensitized; and patients with scores below 0.1 are labeled as lowly sensitized.
We fix the number of incompatible donor-patient pairs to be 100. Donors and patients are randomly sampled from the dataset to form incompatible pairs independently. The overall compatibility is determined by both blood type and HLA compatibility. HLA compatibility is assessed based on the number of mismatches in the A, B, and DR alleles. Specifically, a donor and patient are considered HLA compatible if their level of HLA mismatch is less than 3. The HLA mismatch level is based on UK kidney matching policies, and it can be calculated using the R package transplantr. Again, we only allow cycles of length at most 3.
Figure 3 reports the average selection rates within each subgroup. From Figure 3, group fairness, individual fairness, and our new fairness approach all enhance equitable access to transplant opportunities according to their specific fairness criteria. However, compared to the idealized random graph model discussed in Section 5.1, the price of implementing our new fairness criteria is significantly higher, approximately 8.4%. This discrepancy arises because the PRA scores in the UNOS dataset can vary continuously between 0 and 1, whereas in the previous simulation study in Section 5.1, PRA scores were limited to three discrete values: 0.9, 0.45, and 0.05. Consequently, within each sensitization level, the probability of compatibility with a random donor can differ substantially, increasing the difficulty of balancing selection rates across subgroups. This indicates a potential limitation of our fairness criteria, suggesting the need for a more precise discretization of PRA scores.
Figure 3: Simulation results based on UNOS data. The average selection rates within each subgroup are calculated over 100 data replications. The error bars represent the mean $\pm$ 1 standard deviation of the absolute differences in selection rates.
5.3 Experiment on selection probability prediction
We next conduct a numerical experiment to evaluate the prediction accuracy of selection probability. We assume there is an underlying population consisting of 80% white and 20% non-white incompatible pairs, following the same blood type distributions and sensitization distributions in Section 5.1. We assume the historical data $\{\tilde{v}_{1},...,\tilde{v}_{N_{0}}\}$, current data $\{v_{1},...,v_{N_{1}}\}$, and future data are independently sampled from the population. We fix $N_{0}=200$, $N=100$, $B=1000$, and vary $N_{1}$ in $\{20,40,60,80\}$. For illustration purposes, we focus on our new fairness.
Table 1 presents the numerical results of the selection probability prediction. As $L$ increases, the prediction accuracy improves, evidenced by a lower mean squared error (MSE) and a coverage rate of the prediction interval approaching 95%. This improvement occurs because the size of the unobserved future data, $N-N_{1}$, decreases, which makes the prediction easier. Although these results indicate the method’s validity, there is a bias in the prediction interval due to the finite size of $N_{0}$. When $N_{0}$ is large, the set $\{\tilde{v}_{1},...,\tilde{v}_{N_{0}}\}$ closely approximates the underlying population distribution; while when $N_{0}$ is small, the distribution of $\{\tilde{v}_{1},...,\tilde{v}_{N_{0}}\}$ may differ from the underlying population distribution.
$N_{1}$ 20 40 60 80
MSE 0.087 0.076 0.073 0.059
Coverage 0.986 0.983 0.980 0.981
Width 0.589 0.566 0.541 0.522
Table 1: Accuracy of the predicted selection probability of $v_{1},...,v_{N_{1}}$. The results are averaged over 50 data replications.
Discussion
In this paper, we propose a new fairness criterion that balances selection probabilities within protected groups across each sensitization level. Based on the calibration principle in machine learning, this fairness criterion offers a meaningful and innovative approach in the context of kidney exchange. We propose an efficient solution to implement this criterion and conduct both theoretical and empirical evaluations to analyze the associated price of fairness.
Throughout this paper, we assume that the protected feature $A$ is binary. While it is possible to extend $A$ to a general categorical variable and derive results analogous to Proposition 1, such an extension would alter the upper bound of the number of exchange plans in $\mathcal{F}$ with nonzero selection probability, significantly increasing the computational complexity of the algorithm. We leave it for future research to explore more efficient algorithms to accommodate these potential extensions.
Supplementary material
Section A presents all the mathematical proofs. Section B presents more simulations with failure-ware strategies.
Appendix A Proofs
The proof of Proposition 1 depends on the following lemma.
Lemma S1.
For arbitrary $x_{1},...,x_{n}\in\mathbb{R}^{p}$, any boundary point of the convex hull of the point set $\{x_{1},...,x_{n}\}$ can be represented as $\sum_{i=1}^{n}t_{i}x_{i}$, satisfying that $t_{1},...,t_{n}\geq 0$, $\sum_{i=1}^{n}t_{i}=1$, and the number of nonzero elements in $t_{1},...,t_{n}$ is no greater than $p$.
Proof of Lemma S1.
When $n\leq p$, the result is trial. We only consider the case when $n>p$ in the following proof.
For any point $y\in\mathbb{R}^{p}$ in the convex hull of the point set $\{x_{1},...,x_{n}\}$, we can write $y=\sum_{i=1}^{n}t_{i}x_{i}$, where $t_{1},...,t_{n}\geq 0$ and $\sum_{i=1}^{n}t_{i}=1$. There are different possible choices of $t_{1},...,t_{n}$ satisfying $y=\sum_{i=1}^{n}t_{i}x_{i}$, and here we assume the number of nonzero elements in $t_{1},...,t_{n}$ is minimized. Removing those zero coefficients in $t_{1},...,t_{n}$, we can write $y=\sum_{i=1}^{k}t_{j_{i}}x_{j_{i}}$, where the index set $\{j_{1},...j_{k}\}$ is a subset of $\{1,...,n\}$, $t_{j_{1}},...,t_{j_{k}}>0$ and $\sum_{i=1}^{k}t_{j_{i}}=1$. We only need to show that $k>p$ implies that $y$ must be an interior point of the convex hull of the point set $\{x_{1},...,x_{n}\}$.
We can write
$y-x_{j_{1}}=t_{j_{2}}(x_{j_{2}}-x_{j_{1}})+...+t_{j_{k}}(x_{j_{k}}-x_{j_{1}})=:t_{j_{2}}w_{2}+...t_{j_{k}}w_{k},$
where $t_{j_{2}},...,t_{j_{k}}>0$, $\sum_{i=2}^{k}t_{j_{i}}<1$ and $w_{i}=x_{j_{i}}-x_{j_{1}}$ for $2\leq i\leq k$. We first show that $w_{2},...,w_{k}$ must be linearly independent. Suppose $w_{2},...,w_{k}$ are not linearly independent. Then, there exist $\gamma_{2},...,\gamma_{k}\in\mathbb{R}$ such that $\sum_{i=2}^{k}\gamma_{i}w_{i}=0$ and at least one element in $\{\gamma_{2},...,\gamma_{k}\}$ is nonzero. We consider $\tilde{t}_{j_{i}}=t_{j_{i}}+\delta\gamma_{i}$. Then, $y-x_{j_{1}}=\sum_{i=2}^{k}\tilde{t}_{j_{i}}w_{i}$ and $\sum_{i=2}^{k}\tilde{t}_{j_{i}}<1$ hold for all $\delta\in\mathbb{R}$. By continuity, we can choose $\delta$ such that one element in $\{\tilde{t}_{j_{2}},...\tilde{t}_{j_{k}}\}$ is zero and all other elements are nonnegative. Then
$y-x_{j_{1}}=\tilde{t}_{j_{2}}(x_{j_{2}}-x_{j_{1}})+...+\tilde{t}_{j_{k}}(x_{j_{k}}-x_{j_{1}}),$
which contradicts the minimal choice of $k$. Therefore, we finish the proof that $w_{2},...,w_{k}$ must be linearly independent.
Since $k>p$ and $w_{2},...,w_{k}$ are linearly independent, it is only possible that $k=p+1$. For any $v$ in the unit closed ball in $\mathbb{R}^{p}$, $v$ can be written as a unique linear combination of $w_{2},...,w_{k}$. Due to the compactness of the unit closed ball in $\mathbb{R}^{p}$, there exists $M>0$ such that these absolute values of the linear combination coefficients are less than $M$. Now, we choose
$r=\frac{1}{M}\min\left\{t_{j_{2}},...t_{j_{k}},\frac{1-\sum_{i=2}^{k}t_{j_{i}}}{k-1}\right\}.$
Then, for any $y^{\prime}\in\mathbb{R}^{p}$ satisfying $\|y^{\prime}-y\|_{2}<r$, we can write
$y^{\prime}-y=\lambda_{2}w_{2}+...+\lambda_{k}w_{k},$
where $|\lambda_{i}|<rM$, and thus
$y^{\prime}-x_{j_{1}}=(t_{j_{2}}+\lambda_{2})w_{2}+...+(t_{j_{k}}+\lambda_{k})w_{k}$
where $t_{j_{2}}+\lambda_{2},...,t_{j_{k}}+\lambda_{k}>0$ and $\sum_{i=2}^{k}(t_{j_{i}}+\lambda_{i})<1$. This implies that $y$ is an interior point of the convex hull of the point set $\{x_{1},...,x_{n}\}$. ∎
Proof of Proposition 1.
We encode all the exchange plans in $\mathcal{F}$ based on the vertex inclusion. Let =$\mathcal{F}=\{F_{1},...,F_{S}\}$, where $S=|\mathcal{F}|$, and $F_{i}\in\mathbb{R}^{N}$ where $F_{ij}$ is the binary indicator that the $j$th vertex is included in the exchange plan $F_{i}$. Let $\{w_{1},...w_{S}\}$ be the total utilities of $\{F_{1},...,F_{S}\}$, and let $\{p_{1},...,p_{S}\}$ be the assigned probabilities of $\{F_{1},...,F_{S}\}$. We write $F=(F_{ij})_{S\times N}$, $p=(p_{1},...,p_{S})^{\mathrm{\scriptscriptstyle T}}$, and $w=(w_{1},...,w_{S})^{\mathrm{\scriptscriptstyle T}}$. Then, $q=F^{\mathrm{\scriptscriptstyle T}}p$ is the vector of probability of being matched for each vertex, and $w^{\mathrm{\scriptscriptstyle T}}p$ is the expected utility. Let $\overline{q}_{ij}=|V_{ij}|^{-1}\sum_{v\in V_{ij}}q_{v}$ denote the average selection rate in $V_{ij}$. Then, we can write $(\overline{q}_{11}-\overline{q}_{01},...,\overline{q}_{1M}-\overline{q}_{0M})^{\mathrm{\scriptscriptstyle T}}=Z^{\mathrm{\scriptscriptstyle T}}p$ for some known $S\times M$ matrix $Z$. Let $z_{i}\in\mathbb{R}^{M}$ denote the $i$th row of $Z$, and let $\tilde{z}_{i}=(w_{i}\ z_{i}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{M+1}$. Notice that $Z^{\mathrm{\scriptscriptstyle T}}p$ is a convex combination of $z_{1},...,z_{S}$. We consider $\Omega$ to be the convex hull of $\{\tilde{z}_{1},...,\tilde{z}_{S}\}$. Let $\Delta=[-l_{1},l_{1}]\times...\times[-l_{M},l_{M}]$.
Then, the optimization problem (4) searches over $\Omega\cap(\mathbb{R}\times\Delta)$ and returns an optimum point with the maximum first coordinate. Since $S$ is finite, the convex hull $\Omega$ is compact in $\mathbb{R}^{M+1}$. Thus, $\Omega\cap(\mathbb{R}\times\Delta)$ is compact in $\mathbb{R}^{M+1}$. Furthermore, the optimum point is on the boundary of $\Omega\cap(\mathbb{R}\times\Delta)$, which is also on the boundary of $\Omega$. By Lemma S1, the optimum point can be represented as a convex combination of $\tilde{z}_{1},...,\tilde{z}_{S}$ with at most $M+1$ nonzero coefficients. ∎
The proof of Proposition 2 depends on the following lemma.
Lemma S2.
Consider a random $k$-partite graph $(A_{1},A_{2},...,A_{k})$, where $A_{1},...,A_{k}$ contain $n$ vertices. Each possible edges appears independently with probability no less than $p>0$. Then, the $k$-partite graph has a perfect matching almost surely.
Proof of Lemma S2.
When $k=2$, Erdős and Rényi (1968) gives the probability of perfect matching, which immediately implies Lemma S2. When $k>2$, it can be shown by mathematical induction. We omit the details. ∎
Proof of Proposition 2.
We use $(b_{1},b_{2},r,a)$ to denote the vertex type of an incompatible donor-patient pair with donor blood type $b_{1}$, patient blood type $b_{2}$, patient sensitization level $r$, and patient sensitive group level $a$. Let $\mathbb{T}=\{O,A,B,AB\}\times\{O,A,B,AB\}\times\{r_{1},...,r_{M}\}\times\{0,1\}$. Then, we partition the cycle or relevant subset set $\mathcal{C}$ based on the graph isomorphism with respect to the vertex types in $\mathbb{T}$. Specifically, for $c_{1},c_{2}\in\mathcal{C}$, we view cycle or relevant subset $c_{1}$ and $c_{2}$ of the same type, if there exists a one-to-one mapping $\phi$ from the vertex set in $c_{1}$ to the vertex set in $c_{2}$, satisfying that any $v$ and $\phi(v)$ are of the same type and $(v_{1},v_{2})\in E$ if and only if $(\phi(v_{1}),\phi(v_{2}))\in E$. Therefore, if $c_{1},c_{2}$ are of the same type, then $u(c_{1})=u(c_{2})$. Suppose there are $K$ such types of cycles or relevant subsets in $\mathcal{C}$. Let $U\in\mathbb{R}^{K}$, where $U_{i}=u(c)$ for any cycle or relevant subset $c$ of the $i$th type in $\mathcal{C}$. For arbitrary graph $G$, we encode any exchange allocation as a vector $Z\in\mathbb{R}^{K}$, where $Z_{i}$ denotes the number of cycle or relevant subgraph of the $i$th type in the exchange plan. Let $Opt(G)$ be the solution of optimization problem (7). We claim that the following result holds almost surely as $N\to\infty$, which implies the result in Proposition 2.
  1. 1.
    There exists $Y\in\mathbb{R}^{K}$ such that $Z=\lfloor NY\rfloor\in\mathbb{R}^{K}$ is an achievable exchange allocation in $G$, and $Opt(G)-U^{{\mathrm{\scriptscriptstyle T}}}Z$ is $o(N)$.
  2. 2.
    We can choose $Y\in\mathbb{R}^{K}$ such that $Z=\lfloor NY\rfloor\in\mathbb{R}^{K}$ is an achievable exchange allocation in $G$, where the subgroup $V_{b_{1},b_{2},i,j}$ is prioritized over $V_{b_{1},b_{2},1-i,j}$ for all $(b_{1},b_{2},i,j)\in\mathcal{P}$.
[…]
We can prove the above claims below. Let $(\Omega,\mathcal{F},P)$ denote the probability space. We use $G(N)$ to denote the random graph of size $N$, and we use $G(N;\omega)$ to denote the realized graph of size $N$ for some $\omega\in\Omega$. For fixed $N$, the set $\{Opt(G(N;\omega)):\omega\in\Omega\}$ is bounded, because the number of the edges in $G(N)$ and the edge utilities are finite. Let $S_{N}$ denote the supremum of the set $\{Opt(G(N;\omega)):\omega\in\Omega\}$. Let $Z(N;\omega)\in\mathbb{R}^{K}$ denote the optimal exchange allocation in $G(N;\omega)$. Then, Bolzano–Weierstrass Theorem implies that there exists a subsequence in $\{z(N;\omega):\omega\in\Omega\}$ that converges to $Z_{N}\in\mathbb{R}^{K}$ elementwise, satisfying $U^{{\mathrm{\scriptscriptstyle T}}}Z_{N}=S_{N}$.
For diverging $N$, the sequence $\{S_{N}/N:N\in\mathbb{N}\}$ is bounded, so there exists a subsequence in $\{S_{N}/N:N\in\mathbb{N}\}$ that converges to the supremum of $\{S_{N}/N:N\in\mathbb{N}\}$. Thus, there exists a sequence of increasing numbers $i_{1}<i_{2}<i_{3}<...$ in $\mathbb{N}$, such that $\{S_{i_{k}}/{i_{k}}:k\in\mathbb{N}\}$ converges to the supremum of $\{S_{N}/N:N\in\mathbb{N}\}$. Since the set $\{Z_{i_{k}}/i_{k}:k\in\mathbb{N}\}$ is bounded elementwise, Bolzano–Weierstrass Theorem implies that there exists a subsequence in $\{Z_{i_{k}}/i_{k}:k\in\mathbb{N}\}$ that converges to $Y\in\mathbb{R}^{K}$ elementwise. Based on the construction, we have that $S_{N}/(NU^{{\mathrm{\scriptscriptstyle T}}}Y)$ converges to 1, and thus the supremum of $\{Opt(G(N;\omega))/(NU^{{\mathrm{\scriptscriptstyle T}}}Y):\omega\in\Omega\}$ is no greater than 1.
Next, we show that $Z=\lfloor NY\rfloor$ is an achievable exchange allocation in $G$ almost surely as $N\to\infty$, which will further implies that $Opt(G)-U^{{\mathrm{\scriptscriptstyle T}}}Z$ is $o(N)$. We randomly divide the vertices of $G$ into $K$ disjoint subgraphs based on allocation $Z$, satisfying that the $i$th subgraph contains sufficient vertices to match $Z_{i}$ cycles or relevant subgraphs of the $i$th type. Since some elements in $Z$ can be zero, these corresponding subgraphs can be empty. Lemma S2 guarantees that there is a perfect matching in every subgraph almost surely. Therefore, $Z$ is achievable in random graph $G$ almost surely as $N\to\infty$.
We can manually adjust the elements in $Y$ such that the subgroup $V_{b_{1},b_{2},i,j}$ is prioritized over $V_{b_{1},b_{2},1-i,j}$ for all $(b_{1},b_{2},i,j)\in\mathcal{P}$. Similar to the above argument, $Z=\lfloor NY\rfloor$ is an achievable exchange allocation in $G$ almost surely as $N\to\infty$. We have finished the proof of the above two claims. ∎
The following proofs are based on the expected proportions of subgroups. The observed size of subgroups divided by $N$, should converge to the corresponding expected proportions of subgroups, with high probability. These differences are ignorable in the asymptotic upper bounds of the price of fairness.
Proof of Proposition 3.
For any vertex type $(b_{1},b_{2},r,1)$ and $(b_{1},b_{2},r,0)$, the expected number of these vertices in $G(N)$ is $N(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})$. Suppose that $Np(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})$ of them are expected to be matched in the optimal allocation. Here $p$ depends on $b_{1},b_{2},r$.
If $\mu_{b_{1},b_{2},r,1}/\mu_{b_{1},b_{2},r,0}>\overline{\mu}_{r,1}/\overline{\mu}_{r,0}$, we can only match $np\mu_{b_{1},b_{2},r,0}$ vertices of type $(b_{1},b_{2},r,0)$ and $np\mu_{b_{1},b_{2},r,0}\overline{\mu}_{r,1}/\overline{\mu}_{r,0}$ vertices of type $(b_{1},b_{2},r,1)$ to balance the selection rates of two subgroups. The local relative efficiency loss is no greater than
$1-\frac{np\mu_{b_{1},b_{2},r,0}+np\mu_{b_{1},b_{2},r,0}\overline{\mu}_{r,1}/\overline{\mu}_{r,0}}{np(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})}=\frac{\mu_{b_{1},b_{2},r,1}\overline{\mu}_{r,0}-\mu_{b_{1},b_{2},r,0}\overline{\mu}_{r,1}}{(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})\overline{\mu}_{r,0}}.$
If $\mu_{b_{1},b_{2},r,1}/\mu_{b_{1},b_{2},r,0}\leq\mu_{r,1}/\mu_{r,0}$, we can only match $np\mu_{b_{1},b_{2},r,1}$ vertices of type $(b_{1},b_{2},r,1)$ and $np\mu_{b_{1},b_{2},r,1}\overline{\mu}_{r,0}/\overline{\mu}_{r,1}$ vertices of type $(b_{1},b_{2},r,0)$ to balance the selection rates of two subgroups. The local relative efficiency loss is no greater than
$1-\frac{np\mu_{b_{1},b_{2},r,1}+np\mu_{b_{1},b_{2},r,1}\overline{\mu}_{r,0}/\overline{\mu}_{r,1}}{np(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})}=\frac{\mu_{b_{1},b_{2},r,0}\overline{\mu}_{r,1}-\mu_{b_{1},b_{2},r,1}\overline{\mu}_{r,0}}{(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})\overline{\mu}_{r,1}}.$
Therefore, the overall relative efficiency loss is no greater than
$\max_{b_{1},b_{2},r}\max\left\{\frac{\mu_{b_{1},b_{2},r,1}\overline{\mu}_{r,0}-\mu_{b_{1},b_{2},r,0}\overline{\mu}_{r,1}}{(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})\overline{\mu}_{r,0}},\frac{\mu_{b_{1},b_{2},r,0}\overline{\mu}_{r,1}-\mu_{b_{1},b_{2},r,1}\overline{\mu}_{r,0}}{(\mu_{b_{1},b_{2},r,1}+\mu_{b_{1},b_{2},r,0})\overline{\mu}_{r,1}}\right\}$
almost surely as $n\to\infty$. ∎
Proof of Proposition 4.
Following the notation in Ashlagi and Roth (2011), an X-Y pair has a donor of blood type Y and a patient of blood type X. Without loss of generality, we assume there are more A-B pairs than B-A pairs. By Proposition 5.2 in Ashlagi and Roth (2011), almost surely as $n\to\infty$, there is an optimal allocation such that
  1. (1)
    every pair X-X is matched in a 2-way or a 3-way exchange with other X-X pairs, for X=O,A,B,AB;
  2. (2)
    every B-A pair is matched in a 2-way exchange with A-B pairs;
  3. (2)
    every AB-B pair is matched in a 2-way exchange with B-AB pairs;
  4. (3)
    every AB-A pair is matched in a 2-way exchange with A-AB pairs;
  5. (4)
    every AB-O pair is matched in a 3-way exchange with A-AB pairs and O-A pairs;
  6. (5)
    every A-O pair is matched in a 2-way exchange with O-A pairs
  7. (6)
    every B-O pair is either matched in a 2-way exchange with O-B pairs or in a 3-way exchange with A-B pairs, which are not matched with B-A pairs, and O-A pairs.
From the above optimal allocation, all the X-X, B-A, A-B, AB-B, AB-A, AB-O, A-O, B-O pairs are fully matched, the B-AB, A-AB, O-A, O-B pairs are partially matched, and no O-AB pairs are matched. By Proposition 2, within the blood type pairs that are partially matched, we can give priority to any specific sensitive attribute level $a$ among patients with sensitization level $r$. Based on such an explicit optimal allocation rule, we give an upper bound for the asymptotic price of fairness.
Let $\rho^{-1}$ be the probability that a random patient and a random donor are incompatible. Let $q_{j}$ be the frequency probability of sensitization level $R=r_{j}$. According to the optimal allocation rule, all the X-X, B-A, A-B, AB-B, AB-A, AB-O, A-O, B-O pairs are fully matched, the B-AB, A-AB, O-A, O-B pairs are partially matched, and no O-AB pairs are matched. We partition $V_{i,j}=\{V_{i,j,full}\cup V_{i,j,partial}\cup V_{i,j,none}\}$, where $V_{i,j,full}$ denotes the set of fully matched blood type pairs, $V_{i,r,partial}$ denotes the set of partially matched blood type pairs, and $V_{i,j,none}$ denotes the set of O-AB pairs, within $V_{i,j}$, the set of incompatible pairs of sensitive attribute level $a$ and sensitization level $r$. By Proposition 2, we can arbitrarily arrange the matched pairs in $V_{i,j,partial}$ to achieve fairness. The expected proportion of matched pairs in $\{V_{1,j,partial}\cup V_{0,j,partial}\}$ is equal to $q_{j}$ times the expected proportion of matched pairs in $\cup_{j}\{V_{1,j,partial}\cup V_{0,j,partial}\}$, i.e.,
$\rho q_{j}\overline{\gamma}\sum_{k=0}^{1}\mu_{k}\left\{\mu_{B\mid k}\mu_{AB\mid k}+(\mu_{O\mid k}+\mu_{A\mid k})\mu_{AB\mid k}+(\mu_{A\mid k}+\mu_{AB\mid k})\mu_{O\mid k}+\mu_{O\mid k}\mu_{B\mid k}\right\}.$
Since the expected proportion of $V_{i,j,full}$ is
$\rho\mu_{i}q_{j}\{r_{j}(\mu_{O\mid i}^{2}+\mu_{A\mid i}^{2}+\mu_{B\mid i}^{2}+\mu_{AB\mid i}+\mu_{A\mid i}\mu_{O\mid i}+\mu_{B\mid i}\mu_{O\mid i})+2\mu_{A\mid i}\mu_{B\mid i}\}=\rho q_{j}T_{i}(r_{j})$
and the expected proportion of $V_{i,r}$ is
$\rho q_{j}T_{i}(r_{j})+\rho\mu_{i}q_{j}(\mu_{B\mid i}\mu_{AB\mid i}+\mu_{A\mid i}\mu_{AB\mid i}+\mu_{O\mid i}\mu_{A\mid i}+\mu_{O\mid i}\mu_{B\mid i}+\mu_{O\mid i}\mu_{AB\mid i})=\rho q_{j}S_{i}(r_{j})$
the relative proportion of fully matched pairs within $V_{i,j}$ is $T_{i}(r_{j})/S_{i}(r_{j})$.
We first consider the case when $T_{1}(r_{j})/S_{1}(r_{j})<T_{0}(r_{j})/S_{0}(r_{j})$, under which we should give priority to sensitive attribute level $A=1$ within the sensitization level $R=r$ to balance the selection rates of subgroups.
  1. 1.
    For B-AB pairs, the expected proportion of matched pairs in $\{V_{1,j,partial}\cup V_{0,j,partial}\}$ is
    $\rho q_{j}\overline{\gamma}\sum_{k=0}^{1}\mu_{k}\mu_{B\mid k}\mu_{AB\mid k}$
    while the expected proportion of B-AB pairs in $V_{1,r,partial}$ is $\rho q_{j}\mu_{1}\mu_{B\mid 1}\mu_{AB\mid 1}$. Thus, we can arrange the matched B-AB pairs in $V_{1,j,partial}$ and $V_{0,j,partial}$ such that the difference in expected proportion is as much as
    $\rho q_{j}\min\left\{\sum_{k=0}^{1}\mu_{k}\overline{\gamma}\mu_{B\mid k}\mu_{AB\mid k},2\mu_{1}\mu_{B\mid 1}\mu_{AB\mid 1}-\sum_{k=0}^{1}\mu_{k}\overline{\gamma}\mu_{B\mid k}\mu_{AB\mid k}\right\}.$
  2. 2.
    For A-AB pairs, the expected proportion of matched pairs in $\{V_{1,j,partial}\cup V_{0,j,partial}\}$ is
    $\rho q_{j}\overline{\gamma}\sum_{k=0}^{1}\mu_{k}(\mu_{O\mid k}+\mu_{A\mid k})\mu_{AB\mid k}$
    while the expected proportion of A-AB pairs in $V_{1,r,partial}$ is $\rho q_{j}\mu_{1}\mu_{A\mid 1}\mu_{AB\mid 1}$. Thus, we can arrange the matched A-AB pairs in $V_{1,r,partial}$ and $V_{0,r,partial}$ such that the difference in expected proportion is as much as
    $\rho q_{j}\min\left\{\sum_{k=0}^{1}\mu_{k}\overline{\gamma}(\mu_{O\mid k}+\mu_{A\mid k})\mu_{AB\mid k},2\mu_{1}\mu_{A\mid 1}\mu_{AB\mid 1}-\sum_{k=0}^{1}\mu_{k}\overline{\gamma}(\mu_{O\mid k}+\mu_{A\mid k})\mu_{AB\mid k}\right\}.$
  3. 3.
    For O-A pairs, the expected proportion of matched pairs in $\{V_{1,j,partial}\cup V_{0,j,partial}\}$ is
    $\rho q_{j}\overline{\gamma}\sum_{k=0}^{1}\mu_{k}(\mu_{A\mid k}+\mu_{AB\mid k})\mu_{O\mid k}$
    while the expected proportion of O-A pairs in $V_{1,r,partial}$ is $\rho q_{j}\mu_{1}\mu_{O\mid 1}\mu_{A\mid 1}$. Thus, we can arrange the matched O-A pairs in $V_{1,r,partial}$ and $V_{0,r,partial}$ such that the difference in expected proportion is as much as
    $\rho q_{j}\min\left\{\sum_{k=0}^{1}\mu_{k}\overline{\gamma}(\mu_{A\mid k}+\mu_{AB\mid k})\mu_{O\mid k},2\mu_{1}\mu_{O\mid 1}\mu_{A\mid 1}-\sum_{k=0}^{1}\mu_{k}\overline{\gamma}(\mu_{A\mid k}+\mu_{AB\mid k})\mu_{O\mid k}\right\}.$
  4. 4.
    For O-B pairs, the expected proportion of matched pairs in $\{V_{1,j,partial}\cup V_{0,j,partial}\}$ is
    $\rho q_{j}\overline{\gamma}\sum_{k=0}^{1}\mu_{k}\mu_{O\mid k}\mu_{B\mid k}$
    while the expected proportion of O-B pairs in $V_{1,r,partial}$ is $\rho q_{j}\mu_{1}\mu_{1,B}\mu_{1,AB}$. Thus, we can arrange the matched O-B pairs in $V_{1,r,partial}$ and $V_{0,r,partial}$ such that the difference in expected proportion is as much as
    $\rho q_{j}\min\left\{\sum_{k=0}^{1}\mu_{k}\overline{\gamma}\mu_{O\mid k}\mu_{B\mid k},2\mu_{1}\mu_{O\mid 1}\mu_{B\mid 1}-\sum_{k=0}^{1}\mu_{k}\overline{\gamma}\mu_{O\mid k}\mu_{B\mid k}\right\}.$
The above process leads to a total difference in expected proportion equal to $\rho q_{j}R_{1}$. If $(T_{1}(r_{j})+R_{1})/S_{1}(r_{j})\geq T_{0}(r_{j})/S_{0}(r_{j})$, by continuity, we can always arrange the matched pairs such that the fairness can be perfectly achieved without losing any efficiency. That is, the price of fairness is 00. However, if $(T_{1}(r_{j})+R_{1})/S_{1}(r_{j})<T_{0}(r_{j})/S_{0}(r_{j})$, we can choose to drop some matched pairs in $V_{0,j,full}$ to sacrifice efficiency for fairness. We solve $x$ from the following equation
$\frac{T_{1}(r_{j})+R_{1}}{S_{1}(r_{j})}=\frac{T_{0}(r_{j})-x}{S_{0}(r_{j})}$
and the solution is
$x=\frac{T_{0}(r_{j})S_{1}(r_{j})-T_{1}(r_{j})S_{0}(r_{j})-R_{1}S_{0}(r_{j})}{S_{1}(r_{j})}.$
Thus, the price of fairness within the stratum $R=r_{j}$ is no greater than $x/U(r_{j})$.
We next consider the case when $T_{1}(r_{j})/S_{1}(r_{j})\geq T_{0}(r_{j})/S_{0}(r_{j})$. Similarly, the price of fairness within the level $R=r_{j}$ is no greater than
$\max\left\{\frac{S_{1}T_{0}-S_{0}T_{1}-S_{0}R_{1}}{S_{1}U},\frac{S_{0}T_{1}-S_{1}T_{0}-S_{1}R_{0}}{S_{0}U},0\right\}.$
Therefore, among the whole population, the price of fairness is no greater than
$\max_{r\in\{r_{1},...,r_{M}\}}\max\left\{\frac{S_{1}T_{0}-S_{0}T_{1}-S_{0}R_{1}}{S_{1}U},\frac{S_{0}T_{1}-S_{1}T_{0}-S_{1}R_{0}}{S_{0}U},0\right\}.$
Appendix B More simulations with failure-aware strategies
We repeat the data generating process in Sections 5.1 and 5.2, respectively. With additional vertex and edge uncertainties, let the failure probability $p_{v}$ and $p_{v_{i},v_{j}}$ be independently sampled from a uniform distribution $U(0,0.3)$. We consider the subset-recourse strategy and choose S to be the set of relevant subsets of size $(3,1)$.
Figures S1 and S2 present the average selection rates of different fairness criteria within each subgroup. We do not display the results under weaker fairness constraints of group fairness, individual fairness, and our novel fairness, because they are very close to the results without any fairness constraint. The numerical performance with failure-aware strategies is similar to that without failure-aware strategies.
Figure S1: Simulation results with subset-recourse strategy under random graph models. The average selection rates within each subgroup are calculated over 100 data replications. The error bars represent the mean $\pm$ 1 standard deviation of the absolute differences in selection rates.
Figure S2: Simulation results with subset-recourse strategy based on UNOS data. The average selection rates within each subgroup are calculated over 100 data replications. The error bars represent the mean $\pm$ 1 standard deviation of the absolute differences in selection rates.